# Signatures in TQFT : Asymptotics and Modularity

### v1

@@ 2. $\bullet$
where the inequality in (23) follows from Prop. 4.4(iv), and the inequality (24) follows from $n\leq p_{k}-2$. Plugging this into (22) gives
$|d_{k}|\leq\frac{\pi}{2}\frac{C_{\varepsilon}^{-1}}{\sqrt{p_{k}}}\sum_{n=1}^{\infty}n^{-\frac{3}{2}+\varepsilon}\leq\frac{\widetilde{C}_{\varepsilon}}{\sqrt{p_{k}}}~,$
where $\widetilde{C}_{\varepsilon}$ is a finite constant for $\varepsilon$ small enough ($\varepsilon<\frac{1}{2}$ is enough). Thus $|d_{k}|\rightarrow 0$ as $k\rightarrow\infty$. This proves Lemma 4.8. ∎
Proof of Lemma 4.9.
We assume that $\mu(\theta)=2$ and that $\theta$ satisfies the conditions in Theorem 4.5. Recall $S^{(2)}(q_{k},p_{k})$, defined in (20), is a sum from $n=1$ to $n=p_{k}-2$. We cut it into two parts, as follows. Put
(25) $N(k)=\lfloor 3k/4\rfloor$
and write
$S^{(2)}(q_{k},p_{k})=S^{(2.1)}(q_{k},p_{k})+S^{(2.2)}(q_{k},p_{k})~,$
where $S^{(2.1)}(q_{k},p_{k})$ is the sum from $n=1$ to $n=p_{N(k)+1}-1$ and $S^{(2.2)}(q_{k},p_{k})$ is the sum from $n=p_{N(k)+1}$ to $p_{k}-2$. The following two lemmas provide upper bounds for these two sums.
Lemma 4.10.
We have
$|S^{(2.1)}(q_{k},p_{k})|\leq\frac{p_{N(k)}}{p_{k}}F_{7}(k)$
where $F_{7}(k)={\mathcal{O}}(k^{7})$.
Lemma 4.11.
We have
$|S^{(2.2)}(q_{k},p_{k})|\leq\frac{1}{2}\zeta(2)\frac{p_{k}}{p_{N(k)+1}^{2}}~.$
Assuming these two lemmas for the moment, since $\lim_{m\rightarrow\infty}p_{m}^{1/m}=L>1$ by the Khinchin-Lévy Theorem 4.5(ii), $p_{k}$ grows like $L^{k}$ as $k\rightarrow\infty$. It is not hard to deduce from this that both $S^{(2.1)}(q_{k},p_{k})$ and $S^{(2.2)}(q_{k},p_{k})$ converge to zero.  [^4] Here is how to see this for $S^{(2.1)}(q_{k},p_{k})$. Theorem 4.5(ii) gives that for all $\varepsilon>0$ there exist $m_{0}(\varepsilon)$ so that $L^{1-\varepsilon}\leq p_{m}^{1/m}\leq L^{1+\varepsilon}$ holds for for all $m\geq m_{0}(\varepsilon)$. Thus also $L^{m-m\varepsilon}\leq p_{m}\leq L^{m+m\varepsilon}$ for all $m\geq m_{0}(\varepsilon)$. Applying now the first inequality with $m=k$ and the second one with $m=N(k)=\lfloor 3k/4\rfloor$ and taking $\varepsilon$ small enough, it follows that $p_{N(k)}/p_{k}\leq L^{-\delta k}$ for some $\delta>0$ as soon as $k\geq 2m_{0}(\varepsilon)$. Thus $p_{N(k)}F_{7}(k)/{p_{k}}\leq k^{7}L^{-\delta k}$ for $k$ big enough, proving that the limit as $k\rightarrow\infty$ is zero, as asserted. This proves Lemma 4.9. ∎
It remains to prove Lemmas 4.10 and 4.11. Let us first deal with Lemma 4.11 whose proof is easy.
Proof of Lemma 4.11.
Since $n\geq p_{N(k)+1}$ in this sum, we can bound the $\sin^{2}(n\pi/2p_{k})$ term in the denominator from below by $p_{N(k)+1}^{2}/p_{k}^{2}$. Thus
(26) $|S^{(2.2)}(q_{k},p_{k})|\leq\frac{1}{p_{k}p_{N(k)+1}^{2}}\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{1}{\sin^{2}(q_{k}n\pi/p_{k})}~.$
The sum on the right is bounded from above by
$\sum_{n=1}^{p_{k}-1}\frac{1}{\sin^{2}(q_{k}n\pi/p_{k})}=\sum_{m=1}^{p_{k}-1}\frac{1}{\sin^{2}(m\pi/p_{k})}\leq 2\sum_{m=1}^{\lfloor p_{k}/2\rfloor}\frac{1}{\sin^{2}(m\pi/p_{k})}$
(here we made the substitution $m=q_{k}n\pmod{p_{k}}$ in the first equality). Since $\sin^{2}(m\pi/p_{k})\geq 4m^{2}/p_{k}^{2}$ for $1\leq m\leq\lfloor p_{k}/2\rfloor$, our sum is bounded from above by
[…]
Plugging this into (26) proves Lemma 4.11. ∎
The proof of Lemma 4.10 is more tricky: as we have no lower bound for $n$ in $S^{(2.1)}(q_{k},p_{k})$, both $\sin^{2}$ terms in the denominator can be small at the same time. Still, using an idea of Zaremba (see the proof of [Z, Prop. 2.1]), we can bound the denominator from below as follows. Recall $p_{0}=1$ and the $p_{i}$ are strictly increasing, so every $n<p_{k}$ lies between $p_{i}$ and $p_{i+1}$ for some $i$.
Lemma 4.12.
For $p_{i}\leq n<p_{i+1}$ and $i<k$, we have
(27) $\sin^{2}\Bigl(\frac{\pi n}{2p_{k}}\Bigr)\sin^{2}\Bigl(\frac{q_{k}n\pi}{p_{k}}\Bigr)\geq\frac{4}{p_{k}^{2}(a_{i+1}+2)^{2}}~.$
Proof of Lemma 4.12.
Write
$nq_{k}=x+\ell p_{k},\ \ \ x,\ell\in\mathbb{Z},\ \ \ |x|\leq p_{k}/2~.$
Then using (13) we have
(28) $\sin^{2}\Bigl(\frac{\pi n}{2p_{k}}\Bigr)\sin^{2}\Bigl(\frac{q_{k}n\pi}{p_{k}}\Bigr)=\sin^{2}\Bigl(\frac{\pi n}{2p_{k}}\Bigr)\sin^{2}\Bigl(\frac{x\pi}{p_{k}}\Bigr)\geq\frac{4n^{2}x^{2}}{p_{k}^{4}}~.$
Now using Prop. 4.4(v), we get a lower bound for $|x/p_{k}|$ as follows:
$\Big|\frac{x}{p_{k}}\Big|=\Big|n\frac{q_{k}}{p_{k}}-\ell\Big|\geq\Big|p_{i}\frac{q_{k}}{p_{k}}-q_{i}\Big|\geq\frac{1}{p_{i}+p_{i+1}}$
(since $n<p_{i+1}$.) But then, since $n\geq p_{i}$, it follows that
(29) $\Big|\frac{xn}{p_{k}}\Big|\geq\frac{p_{i}}{p_{i}+p_{i+1}}\geq\frac{1}{a_{i+1}+2}~,$
since (see (16))
(30) $p_{i+1}=a_{i+1}p_{i}+p_{i-1}\leq(a_{i+1}+1)p_{i}~.$
Combining (29) with (28) gives (27). This proves Lemma 4.12. ∎
We are now ready to give the
Proof of Lemma 4.10.
We start with the obvious bound
(31) $|S^{(2.1)}(q_{k},p_{k})|\leq\frac{1}{p_{k}^{3}}\sum_{i=0}^{N(k)}\sum_{n=p_{i}}^{p_{i+1}-1}\frac{1}{\sin^{2}({\pi n}/{2p_{k}})\sin^{2}({q_{k}n\pi}/{p_{k}})}~.$
Applying Lemma 4.12, this is bounded from above by
$\frac{1}{4p_{k}}\sum_{i=0}^{N(k)}\sum_{n=p_{i}}^{p_{i+1}-1}(a_{i+1}+2)^{2}=\frac{1}{4p_{k}}\sum_{i=0}^{N(k)}(p_{i+1}-p_{i})(a_{i+1}+2)^{2}~.$
Applying once more (30), this is bounded from above by
$\frac{1}{4p_{k}}\sum_{i=0}^{N(k)}p_{i}a_{i+1}(a_{i+1}+2)^{2}\leq\frac{p_{N(k)}}{4p_{k}}\sum_{i=0}^{N(k)}a_{i+1}(a_{i+1}+2)^{2}~.$
We now use Khinchin’s Theorem 4.5(i) which tells us that we may assume $a_{i}={\mathcal{O}}(i^{2})$ so that the sum on the right is ${\mathcal{O}}(N(k)^{7})$ hence also ${\mathcal{O}}(k^{7})$. This proves Lemma 4.10. ∎
The proof of Theorem 4.1 is now complete. ∎
@@ 6.2. A modular form related to the genus 2 signature
The following proposition relates certain boundary values of the Eichler integral $G(\tau)$ to the function $\Lambda(\theta)$. Recall that we checked in Section 4 that $\Lambda(\theta)$ converges for diophantine $\theta$. It clearly also converges for $\theta$ rational with even denominator.
@@
For any $\theta$ diophantine or rational with even denominator one has:
@@
The following computation shows that $G(\tau)$ when viewed as a power series in $z^{1/2}=e^{i\pi\tau}$ has convergence radius equal to $1$, and that it also converges for $z^{1/2}=e^{i\pi\theta}$ on the boundary of the disk of convergence when $\theta\in\mathbb{R}$ is diophantine or rational with even denominator, as it converges to $\frac{1}{32}\Lambda(\theta)$ there. Thus the result follows from Abel’s theorem for power series.
@@
For $\theta$ diophantine or rational with even denominator, the function $\Lambda(\theta)$ satisfies the transformation laws $\Lambda(\theta+1)=-\Lambda(\theta)$ and
@@
Beyond proving this conjecture, it would be very interesting to find a direct relation between $\sigma_{2}(\frac{q}{p})$ and the Eichler integral $G$ in the spirit of Section 6.1. It looks also challenging to us to extend these results to higher genus and include marked points in the statements. If $\sigma_{g}(\frac{q}{p})$ indeed grows as $p^{2g-2}$ for any $g\geq 2$ as in Question 5.1, one might hope that its asymptotics is governed by an Eichler integral of a modular form of weight $2g$ as in the genus $2$ case.

### v2

@@ 4. Asymptotic behavior of the signature in genus 22
where the inequality in (25) follows from Prop. 4.3(iv), and the inequality (26) follows from $n\leq p_{k}-2$. Plugging this into (24) gives
$|d_{k}|\leq\frac{\pi C_{\varepsilon}^{-1}}{p_{k}^{\varepsilon}}\sum_{n=1}^{p_{k}}\frac{1}{n}\leq\frac{\pi}{C_{\varepsilon}}\frac{1+\ln(p_{k})}{p_{k}^{\varepsilon}}.$
Thus $|d_{k}|\rightarrow 0$ as $k\rightarrow\infty$. This concludes the proof of Theorem 4.1. ∎
@@ 6.2. A modular form related to the genus 2 signature
The following proposition relates certain boundary values of the Eichler integral $G(\tau)$ to the function $\Lambda(\theta)$. Recall that we checked in Section 4 that $\Lambda(\theta)$ converges for $\theta$ irrational with $\mu(\theta)<3$. It clearly also converges for $\theta$ rational with even denominator.
@@
For any irrational $\theta$ satisfying $\mu(\theta)<3$ or rational $\theta$ with even denominator one has:
@@
The following computation shows that $G(\tau)$ when viewed as a power series in $z^{1/2}=e^{i\pi\tau}$ has convergence radius equal to $1$, and that it also converges for $z^{1/2}=e^{i\pi\theta}$ on the boundary of the disk of convergence when $\theta$ is irrational with $\mu(\theta)<3$ or is rational with even denominator, as it converges to $\frac{1}{32}\Lambda(\theta)$ there. Thus the result follows from Abel’s theorem for power series.
@@
For $\theta$ irrational with $\mu(\theta)<3$ or rational with even denominator, the function $\Lambda(\theta)$ satisfies the transformation laws $\Lambda(\theta+1)=-\Lambda(\theta)$ and
@@
In early February 2026, we were informed by Y. Murakami [Mu] that he has obtained a proof of Conjecture 6.12 starting from our trigonometric formula in Theorem 3.1. A nice feature of his proof is that he shows that the difference between $\frac{1}{p^{2}}\sigma_{2}(q/p)$ and its asymptotic limit is also the boundary value of an Eichler integral, this time of a modular form of weight 2. Still, it would be very interesting to find a direct relation between $\sigma_{2}(\frac{q}{p})$ defined as a sum of signs as in Formula (1) and boundary values of modular forms (or their Eichler integrals) in the spirit of Section 6.1. It looks also challenging to us to extend these results to higher genus and include marked points in the statements. If $\sigma_{g}(\frac{q}{p})$ indeed grows as $p^{2g-2}$ for any $g\geq 2$ as in Question 5.1, one might hope that its asymptotics is governed by an Eichler integral of a modular form of weight $2g$ as in the genus $2$ case.